# Lifting Architectural Constraints of Injective Flows

### v3

@@ 5 Experiments
In implementing the trace estimator, we have to make a number of choices. Briefly, i) we chose to formulate the log-determinant gradient in terms of the encoder rather than decoder as it was more stable in practice, ii) we performed traces in the order $f^{\prime}(x)g^{\prime}(z)$ as this reduces variance, iii) we used a mixture of forward- and backward-mode automatic differentiation as this was compatible with our estimator, and iv) we used orthogonalized Gaussian noise in the trace estimator, to reduce variance. Full justification for these choices is given in appendix D.
@@ 5.1 Tabular Data
Table 1: Free-form injective flows (FIF) are significantly faster than rectangular flows (RF) with superior performance in FID-like metric on 3 out of 4 tabular datasets (Papamakarios et al., 2017). Both methods use $K=1$.
@@
Hyperparameters NLL estimator
& Model (on-/off-manifold) POWER GAS HEPMASS MINIBOONE
We evaluate our method on four of the tabular datasets used by Papamakarios et al. (2017), using the same data splits, and make a comparison to the published rectangular flow results (Caterini et al., 2021), see table 2. We adopt the “FID-like metric” from that work, which computes the Wasserstein-2 distance between the closest Gaussian distributions to the test data and to the data generated by the model. This is a measure of the difference of the means and covariance matrices of the generated and test datasets. We outperform rectangular flows on all datasets except GAS. In addition, we see a speedup in training time of between 1.5 and 6 times between FIF and a rerun of rectangular flows (using the published code) on the same hardware. Full experimental details are in section E.2.
We also perform an ablation to show the effect of the three individual components of our method in table 2: The surrogate of eq. 10, the fix for high curvature solutions in eq. 16 and the use of free-form architectures. Comparing to table 2, we find that the fix to estimate the negative log-likelihood with an off-manifold encoder Jacobian in eq. 7 is crucial for good performance of free-from architectures, as the on-manifold variant in eq. 10 diverges (see section 4.2). This is not the case for the injective architecture based on coupling flows, indicating a stabilizing regularization via the architecture.
@@ 5.2 Comparison to injective flows
We compare FIF against previous injective flows on CelebA images (Liu et al., 2015) in table 3. Our models significantly improves the quality of the generated images in terms of FID. Samples from this model are depicted in fig. 1.
@@ 5.3 Comparison to generative autoencoders
Free-form injective flows (FIF) overcome the architectural constraints of injective flows, and the resulting model can be any autoencoder. We therefore compare FIF to other approaches at training generative autoencoders. This is a general class of bottleneck architectures that encode the training data to a standard normal distribution, so that the decoder can be used as a generator after training.
@@ 6 Conclusion
This paper offers a computationally efficient solution to training normalizing flows on manifolds, which we call the free-form injective flow (FIF). We i) significantly improve an existing estimator for the gradient of the change of variables across dimensions, ii) note that it can be applied to unconstrained architectures, iii) analyze problems with joint manifold and maximum-likelihood training and offer a solution, and iv) implement and test our model on toy, tabular and image datasets. We find that the model is practical and scalable, outperforming comparable injective flows, and showing similar or better performance to other autoencoder generative models on the Pythae benchmark.
Several theoretical and practical questions remain for future work: We identified a problem with jointly learning a manifold and maximum likelihood. We propose a fix in section 4.2 that provides high-quality models, but further investigation is needed for a thorough understanding. Fitting a GMM to the latent space after training improves performance on image data, suggesting that our latent distributions are not perfectly Gaussian. We generally find that architectures with more fully-connected layers in the latent space have a more Gaussian latent distribution, suggesting that larger models suffer less from this problem. We leave potential theoretical or practical improvements to future work.
Acknowledgements
This work is supported by Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy EXC-2181/1 - 390900948 (the Heidelberg STRUCTURES Cluster of Excellence). It is also supported by the Vector Stiftung in the project TRINN (P2019-0092). AR acknowledges funding from the Carl-Zeiss-Stiftung. LZ ackknowledges support by the German Federal Ministery of Education and Research (BMBF) (project EMUNE/031L0293A). We thank the Center for Information Services and High Performance Computing (ZIH) at TU Dresden for its facilities for high throughput calculations. The authors acknowledge support by the state of Baden-Württemberg through bwHPC and the German Research Foundation (DFG) through grant INST 35/1597-1 FUGG.
@@ Appendix D Implementation Details
Considering eqs. 30 and 37, the central component of the surrogate is a trace (estimator). Making use of the cyclic property of the trace, i.e. $\operatorname{tr}(A^{\top}B)=\operatorname{tr}(BA^{\top})$ for any $A,B\in\mathbb{R}^{D\times d}$, we can choose which expansion of the trace to estimate:
@@ D.1 Variance of trace estimator
[…]
@@ D.3.5 Reducing variance when sampling more than 1 Hutchinson sample
A note on our sampling strategy: in practice we sample by taking the $Q$ matrix of the QR decomposition of a $d\times K$ matrix with entries sampled from a standard normal. Since the QR decomposition performs Gram-Schmidt orthogonalization, the $Q$ matrix is uniformly sampled from the group of orthogonal matrices if $Q$ is square (Mezzadri, 2006). The same logic applies to the QR decomposition of non-square matrices, yielding a $d\times K$ matrix made up of the first $K$ columns of a uniformly sampled orthogonal $d\times d$ matrix. Strictly speaking, the QR decomposition is only unique if the $R$ matrix has a positive diagonal, and uniqueness is required for uniform sampling. Let $X=QR$ and define by $D$ the sign of the diagonal of $R$. $D$ is diagonal with $1$ or $-1$ on the diagonal and $D^{2}=\mathbb{I}$. Uniqueness can be achieved by multiplying $Q$ by $D$ from the right and multiplying $R$ by $D$ from the left: $X=QDDR=QR$. The resulting uniformly sampled orthogonal matrix is $QD$, meaning the columns of $Q$ are multiplied by either $1$ or $-1$. In our setting, we have terms of the form $\epsilon_{i}^{\top}A\epsilon_{i}$ where the $\epsilon_{i}$ are the columns of $Q$, so multiplying the columns by $-1$ has no effect on the trace estimate. As a result we opt not to multiply by $D$. Therefore, although we do not sample uniformly from the orthogonal group, the final result is equivalent to sampling uniformly.
@@ E.2 Tabular data
We compare our training times against the published rectangular flow training times for their RNFs-ML ($K=1$) model, as well as rerunning their code on our hardware (a single RTX 2070 card). We find comparable FID-like scores on our rerun (except on GAS where we could not reproduce the score, see section 5), but our hardware is slower, with runs consistently taking at least 15% longer and more than twice as long on MINIBOONE. We find that our model runs in half the time or less of the rectangular flow on the same hardware, except for MINIBOONE (about 2/3 the time).
@@
Training time (minutes)
@@ E.4 Pythae benchmark on generative autoencoders
We compare our method to existing generative autoencoder paradigms using the benchmark from Chadebec et al. (2022). We use the provided open-source pipeline and follow the training setup described by the authors. For MNIST and CIFAR10 this means training for 100 epochs with the Adam optimizer at a starting LR of $10^{-4}$, reserving the last 10k images of the training sets as validation sets. CelebA trains for 50 epochs with a starting LR of $10^{-3}$. All experiments are performed with a batch size of 100 and LR is reduced by half when the loss plateaus for 10 epochs. In accordance with the original benchmark we pick 10 sets of hyper-parameters, compute their validation FID (see fig. 10) and use the model which achieves the best FID on the validation set as the final model. In table 7 we report the FID and IS of this model on the test set. We exclude the VAEGAN from the FID comparison, as the model trains more than double the time required for FIF and goes beyond fitting a transformation of the training data to a standard normal distribution.
As described in section 5, we use the architectures from Chadebec et al. (2022) for the benchmark, which we replicate in tables 8 and 9.
@@ E.5 Compute and dependencies
We build our code upon the following python libraries: PyTorch (Paszke et al., 2019), PyTorch Lightning (Falcon & The PyTorch Lightning team, 2019), Tensorflow (Abadi et al., 2015) for FID score evaluation, Numpy (Harris et al., 2020), Matplotlib (Hunter, 2007) for plotting and Pandas (Wes McKinney, 2010; pandas development team, 2020) for data evaluation.

### v4

@@ 5 Experiments
In implementing the trace estimator, we have to make a number of choices. Briefly, i) we chose to formulate the log-determinant gradient in terms of the encoder rather than decoder as it was more stable in practice, ii) we performed traces in the order $f^{\prime}(x)g^{\prime}(z)$ as this reduces variance (both orderings are valid due to the cyclic property of the trace but since $f^{\prime}(x)g^{\prime}(z)$ is a $d\times d$ matrix whereas $g^{\prime}(z)f^{\prime}(x)$ is $D\times D$, the former is typically easier to estimate), iii) we used a mixture of forward- and backward-mode automatic differentiation as this was compatible with our estimator, and iv) we used orthogonalized Gaussian noise in the trace estimator, to reduce variance. Full justification for these choices is given in appendix D.
@@ 5.1 Tabular Data
Table 1: Free-form injective flows (FIF) are significantly faster than rectangular flows (RF) with superior performance in FID-like metric on 3 out of 4 tabular datasets (Papamakarios et al., 2017). Both methods use $K=1$. The results for RF are taken directly from (Caterini et al., 2021).
@@
We evaluate our method on four of the tabular datasets used by Papamakarios et al. (2017), using the same data splits, and make a comparison to the published rectangular flow results (Caterini et al., 2021), see table 1. We adopt the “FID-like metric” from that work, which computes the Wasserstein-2 distance between the Gaussian distributions with equal mean and covariance as the test data and the data generated by the model. This is a measure of the difference of the means and covariance matrices of the generated and test datasets. We outperform rectangular flows on all datasets except GAS. In addition, we see a speedup in training time of between 1.5 and 6 times between FIF and a rerun of rectangular flows (using the published code) on the same hardware. Full experimental details are in section E.2.
We also show an ablation study in table 6 and table 7 in the appendix disentangling the effect of the three individual components of our method: The surrogate of eq. 10, the fix for high curvature solutions in eq. 15 and the use of free-form architectures. Comparing to table 1, we find that the fix to estimate the negative log-likelihood with an off-manifold encoder Jacobian is crucial for good performance of free-from architectures, as the on-manifold variant diverges (see section 4.2). This is not the case for the injective architecture based on coupling flows, indicating a stabilizing regularization via the architecture.
@@ 5.2 Comparison to injective flows
We compare FIF against previous injective flows on CelebA images (Liu et al., 2015) in table 2. Our models significantly improve the quality of the generated images in terms of the Fréchet inception distance (FID) Heusel et al. (2017) and Inception Score (IS) Salimans et al. (2016). The former compares generated samples to a set of reference samples, by computing the Wasserstein-2 distance between two Gaussian distributions fit to some embedding of the respective sets of samples. The later measures diversity by the entropy of the distribution of class labels in the generated samples, where the class labels are provided by some pre-trained classifier. Samples from this model are depicted in fig. 1.
@@ 5.3 Comparison to generative autoencoders
As free-form injective flows (FIF) do not require any specific architecture, we expand our comparison to the much broader range of generative autoencoders. This is a general class of bottleneck architectures that encode the training data to a standard normal distribution, so that the decoder can be used as a generator after training.
@@ 6 Conclusion
This paper offers a computationally efficient solution to jointly learning a manifold and a distribution on it, which we call the free-form injective flow (FIF). We i) significantly improve an existing estimator for the gradient of the change of variables across dimensions, ii) note that it can be applied to unconstrained architectures, iii) analyze problems with joint manifold and maximum-likelihood training and offer a solution, and iv) implement and test our model on toy, tabular and image datasets. We find that the model is practical and scalable, outperforming comparable injective flows, and showing similar or better performance to other autoencoder generative models on the Pythae benchmark.
Several theoretical and practical questions remain for future work: We identified a previously overlooked problem with jointly learning a manifold and maximum likelihood. We propose a fix in section 4.2 that provides high-quality models, but further investigation is needed for a thorough understanding. Fitting a GMM to the latent space after training improves performance on image data, suggesting that our latent distributions are not perfectly Gaussian. We generally find that architectures with more fully-connected layers in the latent space have a more Gaussian latent distribution, suggesting that larger models suffer less from this problem. We leave potential theoretical or practical improvements to future work.
@@ Appendix D Implementation Details
Considering eqs. 29 and 36, the central component of the surrogate is a trace (estimator). Making use of the cyclic property of the trace, i.e. $\operatorname{tr}(A^{\top}B)=\operatorname{tr}(BA^{\top})$ for any $A,B\in\mathbb{R}^{D\times d}$, we can choose which expansion of the trace to estimate:
@@ D.1 Variance of trace estimator
$E\left[\lVert A^{\top}B\rVert_{F}^{2}\right]=d^{2}D\qquad\textrm{and}\qquad E\left[\lVert BA^{\top}\rVert_{F}^{2}\right]=dD^{2}$ (91)
@@ D.3.5 Reducing variance when sampling more than 1 Hutchinson sample
A note on our sampling strategy: in practice we sample by taking the $Q$ matrix of the QR decomposition of a $d\times K$ matrix with entries sampled from a standard normal. Since the QR decomposition performs Gram-Schmidt orthogonalization, the $Q$ matrix is uniformly sampled from the group of orthogonal matrices if $Q$ is square (Mezzadri, 2007). The same logic applies to the QR decomposition of non-square matrices, yielding a $d\times K$ matrix made up of the first $K$ columns of a uniformly sampled orthogonal $d\times d$ matrix. Strictly speaking, the QR decomposition is only unique if the $R$ matrix has a positive diagonal, and uniqueness is required for uniform sampling. Let $X=QR$ and define by $D$ the sign of the diagonal of $R$. $D$ is diagonal with $1$ or $-1$ on the diagonal and $D^{2}=\mathbb{I}$. Uniqueness can be achieved by multiplying $Q$ by $D$ from the right and multiplying $R$ by $D$ from the left: $X=QDDR=QR$. The resulting uniformly sampled orthogonal matrix is $QD$, meaning the columns of $Q$ are multiplied by either $1$ or $-1$. In our setting, we have terms of the form $\epsilon_{i}^{\top}A\epsilon_{i}$ where the $\epsilon_{i}$ are the columns of $Q$, so multiplying the columns by $-1$ has no effect on the trace estimate. As a result we opt not to multiply by $D$. Therefore, although we do not sample uniformly from the orthogonal group, the final result is equivalent to sampling uniformly.
@@ E.2 Tabular data
We compare our training times against the published rectangular flow training times for their RNFs-ML ($K=1$) model, as well as rerunning their code on our hardware (a single RTX 2070 card). We find comparable FID-like scores on our rerun (except on GAS where we could not reproduce the score, see fig. 3), but our hardware is slower, with runs consistently taking at least 15% longer and more than twice as long on MINIBOONE. We find that our model runs in half the time or less of the rectangular flow on the same hardware, except for MINIBOONE (about 2/3 the time).
@@
Model Training time (minutes)
@@
Hyperparameters NLL estimator
& Model (on-/off-manifold) POWER GAS HEPMASS MINIBOONE
FIF & free-form net off manifold (eq. 15) 0.041 $\pm$ 0.007 0.281 $\pm$ 0.031 0.541 $\pm$ 0.034 0.598 $\pm$ 0.024
Table 7: Reconstruction losses of FIF with a free-form architecture on the POWER, GAS HEPMASS and MINIBOONE datasets. The reconstruction error is always much higher for on-manifold training compared to off-manifold, demonstrating the instability caused by on-manifold NLL evaluation in free-form networks. Note: the large standard deviations in on-manifold runs are typically the result of a single large outlier. We remove the largest outlier where applicable (“On Manifold (outliers removed)” row).
POWER GAS HEPMASS MINIBOONE
On manifold 237 $\pm$ 498 5835 $\pm$ 13006 119 $\pm$ 34 300 $\pm$ 160
On manifold (outliers removed) 14 $\pm$ 27 18 $\pm$ 31 119 $\pm$ 34 229 $\pm$ 23
Off manifold 0.072 $\pm$ 0.002 0.188 $\pm$ 0.012 2.569 $\pm$ 0.098 1.077 $\pm$ 0.011
@@ E.4 Pythae benchmark on generative autoencoders
We compare our method to existing generative autoencoder paradigms using the benchmark from Chadebec et al. (2022). We use the provided open-source pipeline and follow the training setup described by the authors. For MNIST and CIFAR10 this means training for 100 epochs with the Adam optimizer at a starting LR of $10^{-4}$, reserving the last 10k images of the training sets as validation sets. CelebA trains for 50 epochs with a starting LR of $10^{-3}$. All experiments are performed with a batch size of 100 and LR is reduced by half when the loss plateaus for 10 epochs. In accordance with the original benchmark we pick 10 sets of hyper-parameters, compute their validation FID (see fig. 10) and use the model which achieves the best FID on the validation set as the final model. In table 8 we report the FID and IS of this model on the test set. To complement the metrics, we show samples from all models in fig. 11, demonstrating convincing quality. We exclude the VAEGAN from the FID comparison, as the model trains more than double the time required for FIF and goes beyond fitting a transformation of the training data to a standard normal distribution.
As described in fig. 3, we use the architectures from Chadebec et al. (2022) for the benchmark, which we replicate in tables 9 and 10.
@@ E.5 Compute and dependencies
We build our code upon the following python libraries: PyTorch (Paszke et al., 2019), PyTorch Lightning (Falcon & The PyTorch Lightning team, 2019), Tensorflow (Abadi et al., 2015) for FID score evaluation, Numpy (Harris et al., 2020), Matplotlib (Hunter, 2007) for plotting and Pandas (McKinney, 2010; The pandas development team, 2020) for data evaluation.
@@ Appendix F Details on pathology induced by curvature
Adding reconstruction loss alone does not fix this pathology, which we illustrate in fig. 13: The reconstruction loss saturates for small radii, but the best achievable negative log-likelihood (i.e. the entropy of the data) continues to decrease with the radius. Thus, even when increasing $\beta$, the minimal possible value of the total loss is still achieved by a spuriously curved manifold.